A cereal manufacturer is aware that the weight of the product in the box varies slightly from box to box. In fact, considerable historical data has allowed the determination that the weight of the contents is uniformly distributed between 23.75 and 26.25 ounces. What is the probability that a randomly selected box of cereal has product weight less than 24.5 ounces?
Answer #1

Solution :

Given that,

uniformly distribution

a = 23.75

b = 26.25

P(x < c) = (c - a) / (b - a)

P(x < 24.5) = (24.5 - 23.75) / (26.25 - 23.75) = 0.3

Probability = 0.3